One measure used to describe a data set is the range. The range of a data set is given by the difference between the max and the min of the set. The range describes the variability of the data (that is, how much do the numbers vary)
Range = max-min
Find the range of each set of numbers
7, 7, 5, 4, 1, 9, 8, 8, 8, 5, 2

Answer by jim_thompson5910(21667)

To find the range, simply subtract the smallest value from the largest value. Since the largest value is 9 and the smallest is 1, the range is:

Range=max-min=9-1=8

So the range is 8